OpenStudy (anonymous):

Solve: 5(x − 1) − (x + 5) = 2

OpenStudy (anonymous):

Distribute 5(x-1) 5x-5-(x+5)=2 add like terms 4x=2 cause the 5's cancle out. Now isolate x /4 /4 x=2/4 understand? or if u reduce it it would be 1/2

OpenStudy (anonymous):

5x-5-x-5=2 => 4x-10=2 => 4x=12 => x=3